2. a) Distinguish between fiscal policy and monetary policy.
b) Explain three (3) adverse effects a contractionary fiscal policy could have on businesses.
c) Moorgate Ghana Limited has issued 100,000 $1 par equity shares which are at present selling for $4.00 per share. The company has plans to issue rights to purchase one new equity share at a price of $2.50 per share for every three shares.
Required:
i)calculate the theoretical ex rights price of Moogate’s equity shares.
ii)calculate the theoretical value of a Moorgate right, before the shares sell ex rights.
1
Expert's answer
2020-11-19T16:27:16-0500

a)Monetary (or monetary) policy is the management of the money supply in a country. This policy is implemented by the Central Bank. The point is to provide the optimal amount of money in the economy.

Monetary policy tools (methods):

1. Change in the key interest rate.

2. Change in the norm of mandatory reserves.

3. Influence on the exchange rate of the national currency.

Fiscal (fiscal) policy is a government policy that is one of the main methods of state intervention in the economy in order to reduce fluctuations in business cycles and ensure a stable economic system in the short term. The main instruments of fiscal policy are revenues and expenditures of the state budget, i.e. taxes, transfers, and public purchases of goods and services

b)

  • higher taxes;
  • reduction of public procurement - decrease in business income
  • reduction of government subsidies and business subsidies

c)

I) "TERP=\frac{100 000\times4+100 000\times2.5}{100 000+100 000}=3.5"


ii)"TERP=\frac{100 000\times4-100 000\times2.5}{100 000}=1.5"